OpenStudy (gabylovesyou):

@jim_thompson5910 Complete the logical argument by giving a reason for each step ? AB = BC AC = AB + BC AC = AB + AB AC = 2(AB) Given a.________________ b.__________________ c.__________________

OpenStudy (gabylovesyou):



jimthompson5910 (jim_thompson5910):

tell me what you have so far

OpenStudy (gabylovesyou):

i have no idea how to do this

jimthompson5910 (jim_thompson5910):

AC = AB + BC is true for what reason?

OpenStudy (gabylovesyou):

cuz they r even ?

OpenStudy (gabylovesyou):

and if u add it, it creates the whole line

jimthompson5910 (jim_thompson5910):

look through your notes

jimthompson5910 (jim_thompson5910):

it has the word "postulate" in it

OpenStudy (gabylovesyou):

i dont have my notes with me .-.

jimthompson5910 (jim_thompson5910):

they're at school or something?

OpenStudy (gabylovesyou):

segment addition ?

OpenStudy (gabylovesyou):

segment addition postulate ?

OpenStudy (gabylovesyou):

and yes

jimthompson5910 (jim_thompson5910):

good, segment addition postulate

OpenStudy (gabylovesyou):

thats what i put for a. ?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (gabylovesyou):

b. ?

jimthompson5910 (jim_thompson5910):

what changes from AC = AB + BC to AC = AB + AB

OpenStudy (gabylovesyou):

um... idk

jimthompson5910 (jim_thompson5910):

look closely

OpenStudy (gabylovesyou):

if you change the c to a b ?

jimthompson5910 (jim_thompson5910):

so BC turns into AB

jimthompson5910 (jim_thompson5910):

what rule did we just use? and why is this valid?

OpenStudy (gabylovesyou):

because its the same thing.. segment addition ?

jimthompson5910 (jim_thompson5910):

nope not segment addition

OpenStudy (gabylovesyou):

transitive ?

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: we replaced BC with AB

OpenStudy (gabylovesyou):

i forgot the postulate names

OpenStudy (gabylovesyou):

so i have no idea .-.

jimthompson5910 (jim_thompson5910):

hint: this person replaces your teacher when your teacher gets sick

OpenStudy (gabylovesyou):

substitution postulate

jimthompson5910 (jim_thompson5910):

yes, or property

jimthompson5910 (jim_thompson5910):

and that works because AB = BC is given at the top

OpenStudy (gabylovesyou):

and last is multiplication ?